Unit 27 Homework Assignment  Physics141-Summer2009

Unit 27 homework after session 1

• Read chapter 33 in the textbook Understanding Physics.

• Work textbook problems 33-7 and 33-9. Unit 27 homework after session 2

• Work textbook problems 33-47, 33-50 and 33-56.

All Unit 27 homework problems and the unit 27 activity guide and \underline{unit 26 logbook} are the first day of unit 28.